If the perimeter of a rectangular garden is 76 m and the area is 360 square meters, what are the dimensions of the garden
Answer by Cromlix(4381)

Perimeter = 2*length + 2*width
Area = length*width
360 = l * w
length = 360/w
Perimeter = 2*length + 2*width
76 = 2*(360/w) + 2*w
76 = 720/w + 2w = 0
Common denominator = w
76w = 720 + 2w^2
2w^2 - 76w + 720 = 0
Divide by 2
w^2 - 38w + 360 = 0
(w - 18)(w - 20) = 0
Length either:
18m or 20m
Width either
20m or 18m
Hope this helps.
:-)